OpenStudy (anonymous):

For what values of a and b is the line 3 x + y = b tangent to the curve y = a x^{2} when x = 5?

OpenStudy (ash2326):

@jolee 3x+y=b Can you find the slope of this line?

OpenStudy (anonymous):

simple. 3.

OpenStudy (anonymous):

3 is final srry

OpenStudy (ash2326):

think about it

OpenStudy (anonymous):

I tried 3=2ax is this wrong approach? (2ax by taking derivative of y=ax^2

OpenStudy (ash2326):

\[3x+y=b\] taking derivative w.r.t x \[3+\frac{dy}{dx}=0\] \[\frac{dy}{dx}=-3\] which is the slope

OpenStudy (anonymous):

but according to form mx+b, shouldn't slope be 3

OpenStudy (ash2326):

3x+y=b y=-3x+b y=mx+b Comparing both m=-3, b=b

OpenStudy (anonymous):

omg! what a stupid mistake of mine!! thank you

OpenStudy (anonymous):

so a would equal -3/10

OpenStudy (ash2326):

yes, right

OpenStudy (ash2326):

We know that tangent is at point x=5, from this find y coordinate of the point and then use the line equation to find b

OpenStudy (anonymous):

and do u know why we use radians in calc?

OpenStudy (ash2326):

Suppose you have sin x, it's domain is real no.s x=1, 2, 3, these all are radians Using degrees in calculator, you have to convert them first